# CLAPS: Posterior-Aware Conformal Intervals via Last-Layer Laplace

### v2

@@ D.7 Remarks on Misspecification and Diagnostics
computed on a held‐out set (or across calibration folds). These diagnostics instantiate the regimes described by Theorems 4.7 and 4.9:
@@
Consider the following rule with user-chosen small constants $\varepsilon_{r},\varepsilon_{T},\tau_{\rho}>0$:
Select-CLAPS $\iff\mathrm{Median}\{r(x)\}>\varepsilon_{r}\ \text{ and }\ T>\varepsilon_{T}\ \text{ and }\ \rho\leq\tau_{\rho},$
Select-Scale $\iff\mathrm{Median}\{r(x)\}\leq\varepsilon_{r}\ \text{ and }\ \rho>\tau_{\rho},$
where Select-Scale denotes a scale-learning conformal method (e.g., normalized-CP or CQR). Then, under the regimes characterized by Theorems 4.7 and 4.9, the rule aligns with efficiency: it selects CLAPS when posterior shape is informative and selects a scale-learning score when last-layer epistemic variability collapses but heteroscedastic signal is present.
@@
(Contraction regime with heteroscedasticity.) If $T$ is small and $\mathrm{Median}\{r(x)\}\approx 0$, Theorem 4.9 implies $\hat{v}(x)\to\sigma^{2}$ and the posterior-aware score loses its xx–dependent variability. When, in addition, $\rho>\tau_{\rho}$, the residual magnitude correlates with $\sqrt{\hat{v}(x)}$ due to aleatoric structure not captured by the (nearly constant) LLLA variance. Methods that explicitly learn scale (normalized-CP, CQR) can then allocate interval width more efficiently across $x$, yielding narrower sets at the same coverage.
(Remarks on thresholds.) The constants $(\varepsilon_{r},\varepsilon_{T},\tau_{\rho})$ are tuning knobs that define “nontrivial” posterior variance and “substantial” heteroscedastic signal; in practice one may use dataset-specific quantiles (e.g., $\varepsilon_{r}$ as the 10th–20th percentile of $r(x)$ on calibration) rather than fixed numeric cutoffs. The rule is heuristic but consistent with the structural guarantees in Theorem 4.7 and Theorem 4.9. ∎

### v3

@@ D.7 Remarks on Misspecification and Diagnostics
computed on a held‐out set (or across calibration folds). These diagnostics instantiate the regimes described by Theorems 4.7 and 4.10:
@@
Consider the following rule with user-chosen small constants $\varepsilon_{r},\varepsilon_{T},t_{\rho}>0$:
Select-CLAPS $\iff\mathrm{Median}\{r(x)\}>\varepsilon_{r}\ \text{ and }\ T>\varepsilon_{T}\ \text{ and }\ \rho\leq t_{\rho},$
Select-Scale $\iff\mathrm{Median}\{r(x)\}\leq\varepsilon_{r}\ \text{ and }\ \rho>t_{\rho},$
where Select-Scale denotes a scale-learning conformal method (e.g., normalized-CP or CQR). Then, under the regimes characterized by Theorems 4.7 and 4.10, the rule aligns with efficiency: it selects CLAPS when posterior shape is informative and selects a scale-learning score when last-layer epistemic variability collapses but heteroscedastic signal is present.
@@
(Contraction regime with heteroscedasticity.) If $T$ is small and $\mathrm{Median}\{r(x)\}\approx 0$, Theorem 4.10 implies $\hat{v}(x)\to\sigma^{2}$ and the posterior-aware score loses its xx–dependent variability. When, in addition, $\rho>t_{\rho}$, the residual magnitude correlates with $\sqrt{\hat{v}(x)}$ due to aleatoric structure not captured by the (nearly constant) LLLA variance. Methods that explicitly learn scale (normalized-CP, CQR) can then allocate interval width more efficiently across $x$, yielding narrower sets at the same coverage.
(Remarks on thresholds.) The constants $(\varepsilon_{r},\varepsilon_{T},t_{\rho})$ are tuning knobs that define “nontrivial” posterior variance and “substantial” heteroscedastic signal; in practice one may use dataset-specific quantiles (e.g., $\varepsilon_{r}$ as the 10th–20th percentile of $r(x)$ on calibration) rather than fixed numeric cutoffs. The rule is heuristic but consistent with the structural guarantees in Theorem 4.7 and Theorem 4.10. ∎